### v2

# LaF-GRPO: In-Situ Navigation Instruction Generation for the Visually Impaired via GRPO with LLM-as-Follower Reward

Abstract
Navigation instruction generation for visually impaired (VI) individuals (NIG-VI) is critical yet relatively underexplored. This study focuses on generating precise, in-situ, step-by-step navigation instructions that are practically usable for VI users. Specifically, we propose LaF-GRPO (LLM-as-Follower GRPO), where an LLM simulates VI user responses to navigation instructions, thereby providing feedback rewards to guide the post-training of a Vision-Language Model (VLM). This enhances instruction accuracy and usability while reducing costly real-world data collection needs. To address the scarcity of dedicated benchmarks in this field, we introduce NIG4VI, a 27k-sample open-source dataset to facilitate training and evaluation. It comprises diverse navigation scenarios with accurate spatial coordinates, supporting detailed and open-ended in-situ instruction generation. Experiments on NIG4VI demonstrate the effectiveness of LaF-GRPO through quantitative metrics (e.g., Zero-(LaF-GRPO) boosts BLEU 14%; SFT+(LaF-GRPO) METEOR 0.542 vs. GPT-4o 0.323), and qualitative analysis further confirms that our method yields more intuitive and safer instructions.
Codehttps://github.com/YiyiyiZhao/NIG4VI
1 Introduction
Figure 1: A NIG4VI data sample. (a) Bird’s-eye view map with waypoints and corresponding front-view images. (b) A comparison of instructions generated by ASSISTER, the ground truth, and our model (LaF-GRPO).
The Visually Impaired (VI) community, comprising approximately 2.2 billion (WHO2019vision) individuals globally with partial or complete blindness, underscores the significant need for effective assistive technologies. Extensive research in Visually Impaired Assistance (VIA) has emerged to address this need (zhao2024vialmsurveybenchmarkvisually; yuan2025walkvlmaidvisuallyimpairedpeople; gao2025wearable; cai2024navigating). This paper targets a critical and foundational sub-area: Navigation Instruction Generation for the Visually Impaired (NIG-VI). Unlike Navigation Instruction Generation (NIG) for general embodied agents which produces high-level trajectory plans (DBLP:conf/naacl/DouP22; DBLP:conf/acl/GopinathanMAS24; DBLP:conf/eccv/FanLWY24; DBLP:conf/eccv/KongCWSHYL24), NIG-VI is fundamentally human-centered. As illustrated in Figure 1, it demands the generation of in-situ, step-level instructions that must: (1) integrate non-visual sensory cues like auditory/tactile landmarks and surface textures; (2) provide precise directional and distance guidance to compensate for the lack of visual referencing; and (3) dynamically adapt to urban obstacles, all while ensuring user safety within a map-coordinate system.
Early attempts, such as ASSISTER (DBLP:conf/eccv/HuangSZBBO22), laid the initial groundwork in this field but were ultimately constrained by the architectural limitations of BERT-based systems (devlin-etal-2019-bert). The advent of Vision-Language Models (VLMs) has introduced new opportunities due to their multimodal understanding and generation capabilities. Reinforcement Learning (RL) based post-training methods like GRPO (deepseekai2025deepseekr1incentivizingreasoningcapability) further enhance reasoning abilities, enabling VLMs to align with human preferences, i.e., the human-centered guidance demanded by NIG-VI. However, this alignment process requires collecting large-scale human feedback data for fine-tuning, which can be costly and often fails to incorporate the interactive user feedback essential for achieving human-centered guidance.
To bridge this gap, we propose LLM-as-Follower GRPO (LaF-GRPO), the first GRPO-based framework for the NIG-VI task. It features two novel components: (1) an LLM that simulates how VI users respond to navigation instructions by interpreting their likely actions, and (2) a VLM post-training procedure for instruction generation guided by an LLM-as-Follower reward. LaF-GRPO mitigates the need for costly trials with VI users while ensuring instruction usability through a human-in-the-loop navigation simulation. Furthermore, to address the scarcity of VI navigation benchmarks, we introduce NIG4VI—a comprehensive VI navigation instruction benchmark featuring 27k samples for simulation experiments. Fully open-sourced and annotated with granular spatial metadata, NIG4VI enables the generation of detailed, open-ended, in-situ instructions.
Evaluation of LaF-GRPO on the NIG4VI benchmark reveals the following findings: (1) Qwen2.5-VL models under the Zero-(LaF-GRPO) paradigm significantly outperform the standard zero-shot baseline, demonstrating superior performance across a range of metrics. (2) Qwen2.5-VL-7B models trained with SFT+(LaF-GRPO) show leading performance, achieving a METEOR score of 0.542, substantially surpassing powerful proprietary models like GPT-4o. (3) Beyond quantitative improvements, qualitative analysis further confirms that LaF-GRPO generates more human-centered instructions, distinguished by enhanced linguistic variety, more intuitive directional cues, richer environmental details, and crucial safety considerations. In summary, our main contributions are:
$\bullet$ The LaF-GRPO framework, the first to employ GRPO for NIG-VI with a LLM-simulated follower feedback.
$\bullet$ The NIG4VI benchmark, the first open-source comprehensive dataset with precise multi-modal navigation contexts for robust model evaluation for VI navigation.
$\bullet$ Extensive empirical studies across VLMs under various paradigms (Zero-shot, Zero-(LaF-GRPO), SFT, and SFT+(LaF-GRPO)), validating the method’s effectiveness.
2 Related Work
VLMs (DBLP:conf/nips/LiuLWL23a; DBLP:conf/nips/Dai0LTZW0FH23; openai2024gpt4technicalreport; TheC3; geminiteam2024gemini15unlockingmultimodal) have gained attention for combining visual perception with language generation. Refining VLMs with Reinforcement Learning (DBLP:conf/nips/Ouyang0JAWMZASR22) improves alignment with human preferences. Recent success in Group Relative Policy Optimization (GRPO) (deepseekai2025deepseekr1incentivizingreasoningcapability) has led to RL fine-tuned VLMs like AlphaDrive (jiang2025alphadriveunleashingpowervlms), VLM-R1 (shen2025vlmr1stablegeneralizabler1style), Praxis-VLM (hu2025praxisvlmvisiongroundeddecisionmaking), and MedVLM-R1 (pan2025medvlmr1incentivizingmedicalreasoning), broadening their applications.
Visually Impaired Assistance (VIA) is a broad and diverse field (gao2025wearable; cai2024navigating; DBLP:conf/chi/GuanX024; DBLP:conf/chi/LiLKC24). VIA with VLMs is closely related to visual captioning and Visual Question Answering (VQA). VIALM (zhao2024vialmsurveybenchmarkvisually) frames VIA as a VQA task, generating guidance from environment images and user requests. While VIALM emphasizes environment-grounded guidance with tactile information, it is not specifically designed for navigation. WalkVLM (yuan2025walkvlmaidvisuallyimpairedpeople) extends this to dynamic walking assistance and introduces the Walking Awareness Dataset (WAD). Though WalkVLM tackles navigation, its focus remains on video captioning rather than precise orientation and mobility guidance.
There are two main branches for NIG studies: NIG for embodied agents and NIG for the visually impaired. Prior research on Navigation Instruction Generation (NIG) for embodied agents has primarily focused on visual processing while generating trajectory-level instructions. BEVInstructor (DBLP:conf/eccv/FanLWY24) employs a Bird’s-Eye View encoder. SAS (DBLP:conf/acl/GopinathanMAS24) uses semantic knowledge with adversarial reward learning. C-Instructor (DBLP:conf/eccv/KongCWSHYL24) focuses on style-controlled instruction generation. Our proposed approach differs in two significant ways: (1) it prioritizes navigation feedback for VLM fine-tuning; and (2) it generates step-level in-situ instructions.
In the NIG-VI field, ASSISTER (DBLP:conf/eccv/HuangSZBBO22) introduced the UrbanWalk benchmark and developed a navigation assistance model. Our work offers two improvements: (1) we introduce a more detailed evaluation benchmark covering orientation, mobility, scene description, and safety warnings, informed by formative studies from the Human–Computer Interaction field that investigate the needs of VI users (merchant2024generatingcontextuallyrelevantnavigationinstructions; zhao2025lessmorereducingcognitive; DBLP:conf/uist/ChangLG24); and (2) we leverage advanced VLMs within a GRPO framework, leading to more effective instructions.
3 Methodology
Grounded in the Speaker-Follower paradigm (DBLP:conf/nips/FriedHCRAMBSKD18), the rationale for LaF-GRPO is to apply Theory-of-Mind (ToM) principles (DBLP:conf/acl/ZhaoND23). Our LLM-as-Follower simulates a user’s cognitive mapping, generating feedback from their anticipated interpretation.
Figure 2: Method Overview. Top left: An training sample with input, target output, and action interpretation. Bottom left: Action interpreter training using LLaMA-3-8B-Instruct to simulate VI users’ navigation responses. Right: Post-training for VLMs with LaF-GRPO. The LaF reward differentiates outputs when Format and Text Generation rewards are identical.
NIG-VI Task Formulation
In the NIG-VI task, a VLM-based assistant generates in-situ, step-by-step natural language instructions to guide a visually impaired (VI) user along a pre-planned route $P=[p_{1},\dots,p_{K}]$. The route $P$ consists of positional waypoints $p_{i}$ leading to a destination and is generated using the A* algorithm (DBLP:conf/eccv/HuangSZBBO22). At each discrete step $i$ of the navigation, the VLM receives two primary inputs: a front-view camera image $x_{\text{image}}^{(i)}$ and a task description which includes the user’s current pose $x_{\text{pose}}^{(i)}=(x_{\text{loc}}^{(i)},x_{\text{rot}}^{(i)})$ represented by their location $x_{\text{loc}}^{(i)}\in\mathbb{R}^{3}$ and rotation $x_{\text{rot}}^{(i)}\in\mathbb{R}^{3}$ within a global map coordinate system, as well as the next target waypoint $p_{i+1}\in P$. Based on these inputs, the VLM $\pi$ generates a sequence of tokens $y=y^{(i)}=(y_{1}^{(i)},y_{2}^{(i)},\dots,y_{t}^{(i)})$ of token length $t$. The generated instruction $y$ might also include details about the current surroundings captured in $x_{\text{image}}^{(i)}$ and any necessary safety alerts:
$y_{j}\sim\pi_{\theta}(y_{j}^{(i)}|x_{\text{image}}^{(i)},x_{\text{loc}}^{(i)},x_{\text{rot}}^{(i)},p_{i+1},y_{<{j}}^{(i)})$ (1)
where $\theta$ denotes the adjustable model parameters.
The LaF-GRPO Framework
We propose LaF-GRPO to address the challenges of generating navigation instructions that are human-centered and practically usable for VI users, while mitigating the need for costly real-world data collection with VI participants. The overview of LaF-GRPO is illustrated in Figure 2, where the framework comprises two key components: (1) an action interpreter LLM (without a visual encoder to ‘see’) that simulates VI users’ responses to navigation instructions by interpreting how these users would act upon hearing the instructions, and (2) a VLM post-training procedure that generates these instructions with (1)’s feedback.
Action Interpreter. To simulate a VI user’s response to navigation instructions, we use Supervised Fine-tuning (SFT) to train an LLM $\rho_{\varphi}$ to predict potential actions. Prior to its deployment, the interpreter is validated to have a parse accuracy above 98% on a held-out set, ensuring its reliability for generating reward signals. Given VLM-generated instruction tokens $y$, it produces a structured action interpretation $\mathcal{A}$. Formally, we define $\mathcal{A}$ as a structured dictionary containing: (1) a ‘move’ action with associated ‘direction’ (indicated using clock positions (yuan2025walkvlmaidvisuallyimpairedpeople)) and ‘distance’ parameters, and (2) a ‘detailed_hazard_alert’ boolean flag that indicates whether the user perceives warnings about nearby obstacles, as illustrated in Figure 2 Left.
Navigation Instruction Generator. For VI guidance, we use a pre-trained VLM $\pi$ with parameters $\theta$ for in-situ navigation instruction generation. The training of this generator involves two stages: Supervised Fine-tuning (SFT) and Group Relative Policy Optimization (GRPO). Our proposed LaF-GRPO framework enhances standard GRPO (see below) by incorporating a novel LLM-as-Follower reward.
GRPO. The training process of GRPO aims to optimize the policy $\pi_{\theta}$ by maximizing the objective function $\mathcal{J}_{\text{GRPO}}(\theta)$. For a given query $q$, GRPO first samples a batch of $G$ outputs $\{o_{1},o_{2},\ldots,o_{G}\}$ using an older version of the policy, $\pi_{\theta_{\text{old}}}$. The training process of GRPO aims to optimize the policy $\pi_{\theta}$ by maximizing the objective function:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q,\{o_{i}\}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{G}\sum_{i=1}^{G}\mathcal{L}_{i}-\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\right]$ (2)
Here, the term $\mathcal{L}_{i}$ represents the clipped surrogate objective used in PPO (schulman2017proximalpolicyoptimizationalgorithms):
$\mathcal{L}_{i}=\min(w_{i}A_{i},\text{clip}(w_{i},1-\epsilon,1+\epsilon)A_{i})$ (3)
where $w_{i}=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$ is the importance sampling ratio, $A_{i}$ is the estimated advantage for the output $o_{i}$, based on relative rewards of the outputs inside each group only, calculated as $A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\ldots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\ldots,r_{G}\})}$, and $\epsilon$ is a clipping hyperparameter. The second term, $\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})$, regularizes the policy by penalizing divergence from a reference policy $\pi_{\text{ref}}$ with coefficient $\beta$. This regularization stabilizes training by keeping the model close to the original effective policy, preventing it from losing previously learned capabilities.
LaF-GRPO Reward Functions
LaF-GRPO utilizes three reward functions as follows.
Format Reward. To encourage controllable generation, we adopt this binary reward ($r_{\text{format}}\in\{0,1\}$) that evaluates structural compliance with the expected response format. Here, the reward would equal 1 if the output follows the required format pattern $think>\n$<answer>.*?</answer>’ in sequence, and 0 otherwise.
Text Generation Reward. The text generation reward ($r_{meteor}$) is calculated as the METEOR score between the output and the ground-truth reference. METEOR is selected based on its evaluation of semantic overlap, incorporating synonymy and stemming to provide a nuanced assessment.
LLM-as-Follower Reward. The LLM-as-Follower reward ($r_{LaF}$) assesses the navigational quality of generated instructions by comparing their interpreted actions (move direction, distance, and alert flag) with those of a ground-truth reference. The rationale behind this design is that spatial factors, such as directional accuracy ($a_{dir}$) and movement distance precision ($a_{dist}$), play a direct and critical role in determining navigation success. In addition, safety alert flags ($a_{alert}$) serve as supplementary support for VI navigation by indicating potential hazards, though they are not primary determinants of success (giudice2008blind; younis2019smart). Accordingly, we formulate the reward as:
$\begin{split}r_{LaF}=&\,w_{dir}\,\delta(a_{dir},a_{dir}^{ref})+w_{dist}\,\delta(a_{dist},a_{dist}^{ref})\\ &+w_{alert}\,\delta(a_{alert},a_{alert}^{ref})\end{split}$ (4)
$\delta(\cdot)$ denotes an exact match. $r_{LaF}$ is in the range [0, 1].
4 Benchmark: NIG4VI
Benchmark Level # Samples VIA NIG Spatial Acc. Open-ended Open-sourced
R2R (DBLP:conf/cvpr/AndersonWTB0S0G18) High 21k
REVERIE (DBLP:conf/cvpr/QiW0WWSH20) High 10k / 6k
UrbanWalk (DBLP:conf/eccv/HuangSZBBO22) Detailed 2.6k
merchant2024generatingcontextuallyrelevantnavigationinstructions Detailed 48
VIALM (zhao2024vialmsurveybenchmarkvisually) Detailed 200
WAD (yuan2025walkvlmaidvisuallyimpairedpeople) Detailed 12k / 120k
NIG4VI (Ours) Detailed 3k / 24k
- w/o pre-calculation Detailed 1.5k / 12k
- with pre-calculation Detailed 1.5k / 12k
Table 1: Comparison of NIG4VI with existing benchmarks.
We introduce the NIG4VI benchmark to address the scarcity of benchmark resources in this field. Inspired by UrbanWalk, NIG4VI utilizes the open-sourced CARLA Simulator (Dosovitskiy17) to collect samples from diverse scenarios, including varied environments and weather conditions. Pedestrian trajectories are generated using A* algorithm, with precise geospatial coordinates, orientation, frontal-view images, and semantic segmentation images recorded at each step. NIG4VI offers two advantages: (1) its use of a realistic coordinate system facilitates easier transfer to real-world GPS applications, and (2) it enables the cost-effective generation of accurate and extensive data. Table 1 details NIG4VI’s advantages over other datasets.
Dataset Construction. Each question’s input includes the user’s current location/rotation, the next step’s location/orientation, and visual scene data. This information is then structured within a prompt template that includes a detailed task description. The synthesis of the output is a multi-stage process involving both advanced reasoning models and human annotation, as illustrated in Figure 3. Initially, several leading VLMs, specifically GPT-4o, Claude-3.5, and Gemini-2, generate predictions. Following a modality bridging approach, similar to that employed in Vision-R1 (huang2025visionr1incentivizingreasoningcapability), these outputs are processed through DeepSeek-R1 to enhance blindness-oriented spatial guidance and navigability. Crucially, all instructions undergo rigorous human verification. This task is carried out by two annotators, both proficient in English and holding at least an undergraduate-level education, following a similar practice in (zhao2024vialmsurveybenchmarkvisually). The verification involves a two-stage process: first, one annotator performs initial content adjustments, adhering to task requirements. Subsequently, the second annotator independently reviews and verifies this work. Throughout this process, the primary verification criteria include: (1) elimination of visual references (e.g., color-based descriptors), (2) validation of non-visual landmarks, and (3) confirmation of metric precision for mobility-critical parameters.
Figure 3: Dataset instances are first generated using Vision-R1’s modality bridging method (huang2025visionr1incentivizingreasoningcapability) and then reviewed and refined by human annotators.
Dataset Statistics. On average, each town contributes approximately 26.2 navigation routes. The average Euclidean distance between the start and end points of these routes is 113.51 units, with an average of 353.8 steps. After deduplication, the dataset yielded an average of 2,222.7 step-level (image, question) samples per town. It is partitioned into a training set of 1,500 samples from Town01 and a test set. The test set comprises the remaining 613 intra-town samples from Town01, along with all inter-town samples from Town02 (2,579), Town03 (2,260), Town04 (2,316), Town05 (1,935), and Town10 (2,133). Each data sample is available in two versions: ‘with pre-calculation’ and ‘without pre-calculation’. The ‘without pre-calculation’ version requires the VLM to independently calculate navigational parameters (e.g., distance, direction), presenting a greater challenge in guidance generation. Conversely, the ‘with pre-calculation’ version provides the VLM with basic mathematical movement information. The VLM must validate this data and assess the surroundings to generate the navigation instruction.
5 Experimental Settings
Dataset. Experiments utilized the NIG4VI dataset, with Intra-town ($N=613$) and Inter-town ($N=11,223$) test subsets, under ‘with/without pre-calculation’ conditions.
Models. Diverse VLMs were evaluated, falling into two main groups: (1) remote models, such as GPT-4o (openai2024gpt4ocard), Claude-3-5-sonnet-20240620 (Anthropic2024bClaudeSonnet), and Gemini-2.0-flash-thinking-exp-01-21 (GoogleDeepMind2024bGeminiIntroMisc); and (2) smaller, locally runnable VLMs, including DeepSeek-VL-7B (lu2024deepseekvl), LLaVA-v1.6-Mistral-7B (DBLP:conf/cvpr/LiuLLL24), MiniCPM-o-2.6-8B (yao2024minicpm), Intern-VL-2.5-8B (chen2024internvl), and Qwen2.5-VL-3B/7B (Qwen2.5-VL).
Baselines. We compare LaF-GRPO against two primary baseline methods: (1) Zero-shot: Models are applied directly to NIG4VI without prior fine-tuning. (2) Supervised Fine-tuning (SFT): Models are fine-tuned to generate instructions. We implement two variants of LaF-GRPO: (a) Zero-(LaF-GRPO): LaF-GRPO is applied directly to the base model without SFT. (b) SFT+(LaF-GRPO): LaF-GRPO is applied to models that have first undergone SFT.
Evaluation Metrics. Following previous studies in NIG (DBLP:conf/eccv/HuangSZBBO22; DBLP:conf/eccv/FanLWY24; DBLP:conf/eccv/KongCWSHYL24), model performance was evaluated using a suite of widely adopted metrics: BLEU (DBLP:conf/acl/PapineniRWZ02), ROUGE (lin-2004-rouge), METEOR (banerjee-lavie-2005-meteor), and SPICE (anderson2016spicesemanticpropositionalimage). Human studies and LLM-as-Judge evaluations were conducted to assess navigational accuracy and user preference for instruction clarity and helpfulness.
Implementation Details. LaF-GRPO training utilized a single NVIDIA H20 GPU (96 GB of memory). This hardware supports loading an 8B-param LLM (LLaMA-3-8B) and a 3B/7B-param Qwen2.5-VL model for LoRA (hu2022lora) fine-tuning. The reward weights were configured as $(w_{dir},w_{dist},w_{alert})=(0.4,0.4,0.2)$ based on analysis of navigation failure factors, prioritizing spatial parameters over contextual alerts. Training on $3k$ samples took about $15$ hours, with the hyperparameter group size $G$ set to $8$.
6 Results and Discussions
Pre-Cal. Paradigm Model Intra-town ($N$ = 613) Inter-town ($N$ =11,223)
BLEU $\uparrow$ ROUGE $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$ BLEU $\uparrow$ ROUGE $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$
No Zero-Shot DeepSeek-VL-7B 2.179 0.152 0.182 0.116 2.223 0.157 0.196 0.112
MiniCPM-o-8B 2.009 0.145 0.234 0.131 1.969 0.142 0.233 0.129
Intern-VL-8B 1.448 0.150 0.215 0.126 1.517 0.149 0.216 0.120
LLaVA-7B 1.021 0.103 0.201 0.111 1.037 0.107 0.206 0.109
Qwen-VL-7B 3.204 0.202 0.211 0.166 3.128 0.194 0.210 0.157
GPT-4o 1.748 0.169 0.249 0.149 1.617 0.165 0.249 0.142
Claude-3.5 2.803 0.216 0.304 0.211 2.749 0.211 0.301 0.202
Gemin-2 4.105 0.236 0.232 0.232 4.422 0.252 0.238 0.236
Zero-
(LaF-GRPO)
Qwen-VL-3B 3.292 0.230 0.248 0.230 3.972 0.255 0.259 0.244
Qwen-VL-7B 3.272 0.234 0.256 0.222 3.566 0.252 0.260 0.227
SFT Qwen-VL-3B 9.099 0.282 0.496 0.274 8.949 0.284 0.500 0.276
Qwen-VL-7B 9.937 0.291 0.518 0.275 9.709 0.294 0.526 0.281
SFT+
(LaF-GRPO)
Qwen-VL-3B 10.921 0.323 0.528 0.274 10.157 0.309 0.527 0.276
Qwen-VL-7B 10.037 0.284 0.545 0.283 9.002 0.276 0.535 0.278
Yes Zero-Shot DeepSeek-VL-7B 2.517 0.170 0.224 0.161 2.600 0.173 0.237 0.161
MiniCPM-o-8B 2.349 0.166 0.210 0.136 2.517 0.177 0.220 0.144
Intern-VL-8B 1.496 0.132 0.233 0.133 1.517 0.134 0.238 0.132
LLaVA-7B 1.284 0.120 0.222 0.131 1.285 0.121 0.229 0.127
Qwen-VL-7B 2.903 0.188 0.231 0.178 3.080 0.194 0.243 0.180
GPT-4o 2.766 0.204 0.302 0.198 2.967 0.213 0.323 0.211
Claude-3.5 4.124 0.236 0.349 0.257 3.400 0.214 0.326 0.224
Gemin-2 5.132 0.252 0.266 0.269 6.144 0.276 0.283 0.284
Zero-
(LaF-GRPO)
Qwen-VL-3B 3.798 0.249 0.280 0.261 4.584 0.271 0.288 0.274
Qwen-VL-7B 3.678 0.241 0.281 0.229 4.284 0.262 0.286 0.230
SFT Qwen-VL-3B 9.923 0.308 0.512 0.280 10.724 0.318 0.519 0.280
Qwen-VL-7B 9.639 0.270 0.521 0.283 9.710 0.272 0.524 0.287
SFT+
(LaF-GRPO)
Qwen-VL-3B 11.727 0.342 0.541 0.286 10.813 0.333 0.535 0.279
Qwen-VL-7B 10.499 0.285 0.556 0.292 9.232 0.275 0.542 0.288
Table 2: Evaluation results on NIG4VI across Intra-town and Inter-town subsets. Bold values represent the highest score for each metric under a specific setting (with / without pre-calculation), while underlined values indicate the second-highest score. Boxed values highlight the best performing Qwen2.5-VL model within the Zero-Shot and Zero-(LaF-GRPO) categories.
Figure 4: A comparative case study of navigational guidance provided by SFT and SFT+(LaF-GRPO) methods across successive steps. Findings indicate that SFT+(LaF-GRPO) (Ours) generates instructions with greater linguistic variety and more effectively incorporates o’clock directions and specific travel distances.
Pre-Cal.
Reward Types Intra-town ($N$ = 613) Inter-town ($N$ = 11,223)
Format Meteor LLM BLEU $\uparrow$ ROUGE $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$ BLEU $\uparrow$ ROUGE $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$
No $\checkmark$ 10.251 0.318 0.524 0.278 9.401 0.304 0.523 0.279
$\checkmark$ $\checkmark$ 10.912 0.317 0.525 0.279 10.076 0.306 0.521 0.279
$\checkmark$ $\checkmark$ $\checkmark$ 10.921 0.323 0.528 0.274 10.157 0.309 0.527 0.276
Yes $\checkmark$ 11.269 0.337 0.538 0.292 10.217 0.328 0.530 0.282
$\checkmark$ $\checkmark$ 11.602 0.339 0.539 0.284 10.753 0.331 0.531 0.280
$\checkmark$ $\checkmark$ $\checkmark$ 11.727 0.342 0.541 0.286 10.813 0.333 0.535 0.280
Table 3: Ablation study results for the Qwen2.5-VL-3B model on the NIG4VI dataset with different reward functions.
Pre-Cal.
Model Intra-town ($N$ = 613) Inter-town ($N$ = 11,223)
BLEU $\uparrow$ ROUGE $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$ BLEU $\uparrow$ Rouge $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$
No 7B-format+meteor+LLM (1k) 9.401 0.283 0.529 0.274 8.963 0.281 0.530 0.275
7B-format+meteor+LLM (2k) 9.657 0.280 0.539 0.276 9.001 0.276 0.535 0.274
7B-format+meteor+LLM (3k) 10.037 0.284 0.545 0.283 9.002 0.276 0.535 0.278
Yes 7B-format+meteor+LLM (1k) 10.265 0.279 0.543 0.286 9.463 0.271 0.540 0.285
7B-format+meteor+LLM (2k) 10.136 0.284 0.550 0.292 9.245 0.276 0.541 0.284
7B-format+meteor+LLM (3k) 10.499 0.285 0.556 0.292 9.232 0.275 0.542 0.288
Table 4: Ablation study results for the Qwen2.5-VL-7B model on the NIG4VI dataset with varying training sample sizes.
Main Results
Table 2 summarizes model performance on NIG4VI, categorized by pre-calculation and training paradigms, and evaluated on intra-town and inter-town subsets. Comparing LaF-GRPO with the baselines reveals: (1) Zero-Shot vs. Zero-(LaF-GRPO): Zero-(LaF-GRPO) significantly enhances the Zero-Shot performance of VLMs, validating the effectiveness of LaF-GRPO. While the Zero-(LaF-GRPO) results suggest that increased model size (from 3B to 7B) does not necessarily guarantee improved performance across all metrics, it is noteworthy that for METEOR evaluations, specifically in intra-town scenarios, the 7B model achieved the highest scores (i.e., $0.256$ and $0.281$). This outcome may be attributable to the use of METEOR as a text generation reward during training and to the potentially more refined tuning applied to the 7B models. (2) SFT & SFT+(LaF-GRPO): SFT and SFT+(LaF-GRPO) yield significantly superior performance compared to Zero-Shot and Zero-(LaF-GRPO) models across all metrics and subsets, affirming the efficacy of fine-tuning. The SFT+(LaF-GRPO) approach further enhances performance beyond SFT. Moreover, under the SFT+(LaF-GRPO) paradigm, Qwen-VL-3B consistently achieves the highest BLEU and ROUGE scores, while Qwen-VL-7B excels in METEOR and SPICE. This performance pattern is observed for both intra-town and inter-town subsets and holds true regardless of pre-calculation. This may be attributable to 7B models demonstrating enhanced linguistic diversity in their outputs relative to 3B models. (3) Additional Observations: Scores improved with pre-calculation, which reduced computational difficulty, whereas the more challenging ‘w/o pre-calculation’ setting caused more failures. Intra-town scores were higher than inter-town, likely due to a closer data distribution with the training set. We also observed that LaF-GRPO generated significantly more concise, user-friendly instructions (e.g., LaF-GRPO-7B:  34.1 tokens vs. GPT-4o:  117.9 tokens).
Ablation Studies
Reward Types. Table 3 presents an ablation study investigating the impact of different reward types during SFT+(LaF-GRPO) training with the Qwen-VL-3B model. LaF-GRPO, incorporating the LLM-as-Follower reward, consistently achieves the highest BLEU, ROUGE, and METEOR scores. This trend holds true across both intra-town and inter-town evaluations, with or without pre-calculation. This underscores the significant benefit of the LaF reward.
Training Sample Sizes. Table 4 presents an ablation study on the Qwen2.5-VL-7B model trained with SFT+(LaF-GRPO), illustrating the effect of varying training sample sizes ($1k$, $2k$, and $3k$). For comprehensive metrics such as METEOR and SPICE, performance generally scales with the volume of training data. Across the majority of evaluated conditions, scaling up to $3k$ samples typically yields the optimal or near-optimal scores. Nevertheless, training with $2k$ samples also achieves comparable METEOR scores, indicating training data efficiency at this sample size.
LaF-GRPO vs. Standard GRPO. We conducted LLM-as-Judge and human evaluations on the inter-town subset. The LLM-as-Judge evaluation with GPT-4o showed LaF-GRPO’s superior navigational accuracy (68.1% vs. 67.3%) and a greater preference rate for instruction clarity and helpfulness (58.3% vs. 41.7%). These findings were corroborated by a human preference study on 100 instruction pairs, evaluated by two trained graduate students acting as VI user simulators. To isolate instruction quality, the evaluators were instructed to navigate relying solely on text and ignoring visual cues. Their assessments achieved substantial inter-rater agreement (Cohen’s $\kappa$= 0.83), and the results confirmed a strong preference for LaF-GRPO instructions (76% vs. 24%) and its higher navigational accuracy (79.0% vs. 77.5%). We identify two main advantages of LaF-GRPO over standard GRPO, which utilizes only format and text generation rewards: (1) Navigational Accuracy: LaF-GRPO provides more precise movement and orientation guidance. (2) Instruction Clarity: The action interpreter component encourages VLMs to generate instructions that are clearer, more structured, and more comprehensible.
Case Study
Figure 4 provides a qualitative comparison of SFT+(LaF-GRPO) against the SFT baseline. Notably, SFT+(LaF-GRPO) generates instructions with greater linguistic variety and more intuitive directional cues. For instance, in Step 2, SFT+(LaF-GRPO) employs an o’clock direction (“Turn slightly right toward 1 o’clock”) and a relatable distance (”two small steps”), contrasting with SFT’s numerical bearing (”150 degrees”). It can yield guidance that is more naturally understood by VI users. Furthermore, SFT+(LaF-GRPO), leveraging its internal reasoning process (i.e. the <think>...</think> blocks), frequently incorporates more environmental details and safety considerations. For example, its instruction for Step 4 (”Step forward 0.5 meters; …use your cane near the left … and listen for traffic”) also emphasizes immediate safety interactions.
Limitations and Future Work
This study’s reliance on simulation and proxy users, while necessary for large-scale, controlled, and safe initial testing, introduces limitations. Future work could explore real-world dataset collection and direct engagement with VI users.
7 Conclusion
This study addresses navigation instruction generation for the visually impaired individuals (NIG-VI). We constructed the NIG4VI benchmark. Following this, we developed LaF-GRPO, a novel training paradigm for VLMs that incorporates an LLM-as-Follower reward. Experimental evaluations established LaF-GRPO’s superiority over baselines and standard GRPO, with qualitative analysis confirming the practicality of the generated instructions in real-world scenarios. We hope our work and the benchmark inspire the development of more effective aids for the visually impaired.
Acknowledgments
This work was supported by the Innovation and Technology Fund (Project No. PRP/047/22FX), a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU/25200821), and PolyU Internal Fund from RC-DSAI (Project No. 1-CE1E). We also thank the reviewers for their constructive comments.